Consider the example in class, where we have data on workers Earning, Schooling years and Gender. And the model is specified as: Earning = β0 + β1Schooling + β2G + u where the model satisfies the 4 assumptions. And G=1 if the worker is female. (1) Interpret the effect of schooling years on male worker’s earning. (Hint: E(Earning|G = 0)) (2) Interpret the effect of schooling years on female worker’s earning.
Answer #1

Earning = β0 + β1Schooling + β2G + u

(1)

For male worker’s G = 0,

then,

Earning = β0 + β1Schooling + u

With the increase in 1 year of schooling, the male worker’s earning will increase by β1.

Without schooling (schooling = 0), the male worker’s earning is β0.

(2)

For male worker’s G = 1,

then,

Earning = β0 + β2 + β1Schooling + u

With the increase in 1 year of schooling, the female worker’s earning will increase by β1.

Without schooling (schooling = 0), the female worker’s earning is β0 + β2.